Controls Meeting 23-Feb-2012
- Announcements (concerning solenoid, from Tim)
- Chimney connections in Sept/OCT
- Controls and low power tests Oct/Nov
- Cold Nov/Dec
- Full Power Jan/Feb
- Minutes from previous meeting 9-Feb-2012
- Counting house installation, power, UPS power, fibers, Tagger bldg, etc. - Hovanes, Mark, Elliott
- Equipment orders - Mark, Hovanes
- Solenoid_Controls_Redesign - Mark, Josh, Scot, Yi, Elliott
- SOE circuit, installation and programming
- QD, shunt replacement, ground fault detector, racks, UPS power to racks
- NI PXI device testing
- Experiment Control Supervisor - Elliott
- EPICS - Hovanes, Nerses
- archiver, alarm system, scalers, etc.
- Test stand - Hovanes, Josh
- Target, goniometer, other new systems - Hovanes, Elliott
- PLC programming - Josh
- talking to TCP via EWEB (or EN2T?)
- program redesign, new features, new computer
- Elogs and databases - Elliott
- Midas elog
- Operator elog
- inventory database - CED, CLAS, IRMIS?
- Mantis task tracker
9:30 AM Thurs 23-Feb-2012 CC L207
9:30am 8-Mar-2012 CC L207
Present: Elliott W, Mark S, Josh B.
Tim's schedule has cryo shakedown starting by Nov, and cold powered operation in Jan.
Installation is proceeding on the UPS and other items. Giles et al. determined that supports are not needed under the floor panels. Only one side of the Rack Room has adequate cooling return, this will have to be added to the other side when we upgrade and install more computers.
Mark will try one more time to hear from the Danfysik engineer before ordering an isolation amp to directly read out the ZFCT.
Solenoid Controls Redesign
- We agreed to implement Mark's proposed scheme for hooking up the SOE module to the trip system. Mark identified a number of trip circuits we need to connect to. Josh says programming the SOE is simple and straightforward.
- Concerning the ground fault detector, Mark will measure the dump resistor to find the center, then have someone weld a tab there.
- Yi got a loaner NI PCX device and is currently testing it with Josh's help. He was able to create an RT program that reads out some modules, and now will test connection to the PLC.
- Mark proposes we read out the ZFCT with an isolation amplifier as a replacement for the shunt. This needs to be tested with the power supply connected to line power and cooling water. Elliott will get the schedule for this from Tim.
Experiment Control Supervisor
Vardan is working out how to integrate the EPICS slow control and DAQ run control under the CSS framework. This work is on hold until the DAQ group gets the CODA3 alpha release out.
- We likely will use the Accelerator archiver once they connect it to the CSS framework.
- Hovanes is evaluating the CSS alarm system.
- We talked with Accelerator about using their updated backup/restore system.
- Nerses developed EPICS drivers for JLab scaler readout.
- Nerses is turning his attention to HV and LV control.
Chris Keith now plans to use one Lakeshore temperature monitor and one Lakeshore temperature controller, due to the increase in the number of channels. One concern was how to talk to the new controllers that do not have RS232 connections. Josh found an RTI module (490NBX) that converts from Ethernet/IP to raw TCP that should do the trick.
Josh may look at raw TCP programming using "socket services" provided by the EWEB module, although the RTI 490NBX probably means this is not needed. The module costs around $650 for 10 channels. So far we know of one channel needed, but more will likely appear.
Accelerator may upgrade their elog to meet requirements from the four halls. Graham is contact person for the halls. We shall see what comes up...
- We need some sort of base inventory fairly soon.
- We are looking at three systems: CED from Accelerator, a system from CLAS, and IRMIS3 from the EPICS community.
- We will be testing CED and the CLAS system next week.
- It is not clear when IRMIS3 will be released.
- We might start with one and migrate to another later.
- An enhanced inventory system may replace Midas as a maintenance log, not clear yet.